# Gold-medalist Performance in Solving Olympiad Geometry with AlphaGeometry2

### v1

@@ Abstract
We present AlphaGeometry2, a significantly improved version of AlphaGeometry introduced in Trinh et al. (2024), which has now surpassed an average gold medalist in solving Olympiad geometry problems. To achieve this, we first extend the original AlphaGeometry language to tackle harder problems involving movements of objects, and problems containing linear equations of angles, ratios, and distances. This, together with other additions, has markedly improved the coverage rate of the AlphaGeometry language on International Math Olympiads (IMO) 2000-2024 geometry problems from 66% to 88%. The search process of AlphaGeometry2 has also been greatly improved through the use of Gemini architecture for better language modeling, and a novel knowledge-sharing mechanism that combines multiple search trees. Together with further enhancements to the symbolic engine and synthetic data generation, we have significantly boosted the overall solving rate of AlphaGeometry2 to 84% for all geometry problems over the last 25 years, compared to 54% previously. AlphaGeometry2 was also part of the system that achieved silver-medal standard at IMO 2024 https://dpmd.ai/imo-silver. Last but not least, we report progress towards using AlphaGeometry2 as a part of a fully automated system that reliably solves geometry problems directly from natural language input.
@@ 1 Introduction
The International Mathematical Olympiad (IMO) is a prestigious mathematics competition for high school students worldwide. IMO problems are known for their difficulty, and solving them requires deep understanding of mathematical concepts and the ability to apply them creatively. Geometry, one of the four IMO categories, is the most uniform across problems, hence the most approachable. It is also well suited for basic reasoning research.
@@
There has been two main approaches for automatically solving geometry problems. One is bashing the problems algebraically with Wu’s method Chou (1985); Wu (2008), Area method Chou et al. (1993, 1994), or Gröbner bases Kapur (1986a, b), and second approach relies in synthetic techniques such as Deduction database Chou et al. (2000), or Full angle method Chou et al. (1996). We focus on the latter as a more human-like approach suitable for transferring the research knowledge to other domains. In our previous work Trinh et al. (2024), we introduced AlphaGeometry (AG1), a neuro-symbolic system that demonstrated a significant step towards mastering this domain, achieving a 54% solve rate on all 2000-2024 IMO geometry problems. AG1 combines a language model (LM) with a symbolic engine to effectively tackle these challenging problems.
Despite its success, AG1 exhibited limitations in several key areas. Its performance was constrained by the scope of its domain-specific language, the efficiency of the symbolic engine, and the capacity of the initial language model. As a result, when considering all the recent IMO geometry problems from the year 2000 until now, AG1 can only achieve a solving rate of 54%.
@@
This paper introduces AlphaGeometry2 (AG2), a substantial upgrade that addresses these limitations and significantly enhances performance. AG2 leverages a more powerful Gemini-based language model trained on a larger and more diverse dataset. We also introduce a significantly faster and more robust symbolic engine, incorporating optimizations such as a reduced rule set and enhanced handling of double points. Furthermore, we expand the domain language to encompass a wider range of geometric concepts, including locus theorems and linear equations. To further improve performance, we develop a novel search algorithm that explores a broader range of auxiliary construction strategies, and employs a knowledge-sharing mechanism to expand and accelerate the search process. Finally, we make progress towards building a fully automated and reliable system that solves geometry problems in natural language. To do this we utilize Gemini Team Gemini (2024) to translate problems from natural language into the AlphaGeometry language and implement a new automated diagram generation algorithm.
@@
These enhancements culminate in a substantial improvement in performance: AG2 achieves an impressive 84% solve rate on all 2000-2024 IMO geometry problems, demonstrating a significant leap forward in AI’s ability to tackle challenging mathematical reasoning tasks, and surpassing an average IMO gold medalist.
AlphaGeometry2 Key Improvements:
@@
  • Expanded Domain Language: Covering locus-type theorems, linear equations, and non-constructive problem statements.
@@
  • Stronger and faster Symbolic Engine: Optimized rule set, added handling of double points, and a faster implementation in C++.
  • Advanced Novel Search Algorithm: Utilizing multiple search trees with knowledge sharing.
@@
  • Enhanced Language Model: Leveraging the Gemini architecture, trained on a larger and more diverse dataset.
@@ 3 Automated problem formalization and diagram generation
Automating this process, called formalization, is an active area of research (see for example, Szegedy (2020); Wu et al. (2022); Jiang et al. (2023); Poiroux et al. (2024)). It is a significantly more complicated problem compared to a translation between human languages. While translation aims to preserve meaning, formalization frequently requires re-formulating the original problem into an alternative form, and sometimes disambiguating the nuances in the original problem statement. Automated formalization (auto-formalization), therefore, demands significant background knowledge and problem-solving skills on its own. Given that recently foundation models started demonstrating such capabilities, we use one such model, Gemini Team Gemini (2024), to automate the problem formalization for AlphaGeometry. We start by manually translating several dozens of geometry problems into the AG language. Then we use these examples to write a few-shot prompt asking Gemini to translate a given geometry problem from natural language into the AG language. We query Gemini five times with this prompt followed by another Gemini call asking to combine these results into one final answer. With this approach we are able to formalize 30 out of 39 formalizable IMO 2000-2024 geometry problems. For easier geometry problems, it is very consistent and makes almost no mistakes.
@@
Another manual part of our pipeline was diagram generation. In AG1, each point is defined by at most two basic predicates recalled in Table 1, the problem is therefore defined constructively and diagrams can be generated automatically. In AG2, we allow one or multiple points being defined simultaneously by an arbitrary number of predicates, allowing us to also cover non-constructive problems. Consider a non-constructive problem statement, “Let $ABC$ be a triangle with incenter $I$, such that $IA=2IB$”, here point $I$ is not only defined as an incenter, i.e. the intersection of two internal bisectors, but also defined by a third predicate $IA=2IB$ and there is no general strategy to construct such four points. Since AG2 covers non-constructive problems, diagram construction becomes a non-trivial part of the pipeline and generally requires human intervention. Similar to Krueger et al. (2021), we propose the following algorithm to automatically generate diagrams given non-constructive problem specifications:
Let $\bar{x}\in\mathbb{R}^{2n}$ be a vector representing all coordinates of all points. We encode every constraint $c$ in the diagram, including the goal, as $f_{c}(\bar{x})=0$ with a nonlinear function $f_{c}$. We numerically search for a suitable $\bar{x}$ in two steps. First, we run the ADAM gradient descent optimization on the mean-squared error loss $\sum_{c\in C}f_{c}(\bar{x})$ where $C$ is the set of all the constraints, together with a non-degeneracy loss. For every two points $A$, $B$, we add the loss of the form $1/{(|AB|^{2}+\epsilon)}$, and an $L_{2}$ normalization on all points to prevent their value from becoming too large. After the loss in the ADAM optimization meets a certain threshold, we stop caring about the non-degeneracy, and switch from a gradient descent optimization to the Gauss-Newton-Levenberg method to look for a numerical solution of a combined under- and over-determined system of nonlinear equations.
@@
This two-stage optimization method builds upon the methodology introduced in Krueger et al. (2021). While the first stage remains unchanged, we incorporate a novel second stage. This addition addresses the practical limitations encountered when tuning the gradient descent optimization in the original method, where achieving a consistently satisfactory error margin proved challenging.
@@
We benchmark this method on 44 IMO problems formalized in AG language (see Figure 8) and are able to find diagrams for 41 of them. We run the two-stage convergence procedure in multiple parallel processes, and in a loop which restarts and generates another random initial configuration after a failure. This way, 40 / 44 problems got their diagram generated within 1 hour using approximately 40 processes for each problem (many problems got their diagram within seconds, on the first try). For the remaining 4 problems, we run the same procedure longer and with more parallelization. This way, we also obtained a diagram for IMO-2011-6 after 400 minutes using $3$ processes.

### v2

@@ Abstract
We present AlphaGeometry2, a significantly improved version of AlphaGeometry introduced in Trinh et al. (2024), which has now surpassed an average gold medalist in solving Olympiad geometry problems. To achieve this, we first extend the original AlphaGeometry language to tackle harder problems involving movements of objects, and problems containing linear equations of angles, ratios, and distances. This, together with support for non-constructive problems, has markedly improved the coverage rate of the AlphaGeometry language on International Math Olympiads (IMO) 2000-2024 geometry problems from 66% to 88%. The search process of AlphaGeometry2 has also been greatly improved through the use of Gemini architecture for better language modeling, and a novel knowledge-sharing mechanism that enables effective communication between search trees. Together with further enhancements to the symbolic engine and synthetic data generation, we have significantly boosted the overall solving rate of AlphaGeometry2 to 84% for all geometry problems over the last 25 years, compared to 54% previously. AlphaGeometry2 was also part of the system that achieved silver-medal standard at IMO 2024 https://dpmd.ai/imo-silver. Last but not least, we report progress towards using AlphaGeometry2 as a part of a fully automated system that reliably solves geometry problems directly from natural language input.
@@ 1 Introduction
Advancing reasoning capabilities of artificial intelligence (AI) systems, especially in tackling complex mathematical problems, has been a decade-long holy grail of AI research. Even with the rapid developments of large language models, most systems until now still struggle with fundamental understanding of geometric objects (e.g., points, lines, circles) and their interactions. This had motivated our previous work, AlphaGeometry Trinh et al. (2024), a neuro-symbolic system capable of solving Euclidean geometry problems at the Olympiad level. Specifically, AlphaGeometry (AG1) combines the power of a neural language model, that is creative in suggesting insights for tackling challenging problems, with a symbolic engine, that uses a formal language to describe geometric constructs and deduction rules to reliably derive new properties. Trained on 100M synthetically generated examples of diverse levels of complexity, AG1 was able to solve geometry problems at the level of the International Mathematical Olympiad (IMO), a prestigious competition for the world’s brightest high-school mathematicians.
@@
Broadly speaking, there are two main approaches for automatically solving geometry problems. The first group is about bashing the geometry problems algebraically with Wu’s method Chou (1985); Wu (2008), Area method Chou et al. (1993, 1994), or Gröbner bases Kapur (1986a, b). The second group, on the other hand, relies on synthetic techniques, such as Deduction database Chou et al. (2000), or Full angle method Chou et al. (1996). AG1 uses the latter, as a more human-like approach suitable for transferring the research knowledge to other domains. The neuro-symbolic system that AG1 employed has demonstrated a significant step towards mastering the domain of Euclidean geometry. However, despite its success, AG1 exhibited limitations in several key areas. Its performance was constrained by the scope of its domain-specific language, the efficiency of the symbolic engine, and the capacity of the initial language model. As a result, when considering all the recent IMO geometry problems from the year 2000 until now, AG1 can only achieve a solving rate of 54%.
@@
This paper introduces AlphaGeometry2 (AG2), a substantial upgrade that addresses the aforementioned limitations and significantly enhances performance. First, we expand the domain language to encompass a wider range of geometric concepts, including locus theorems with moving constructs and linear equations of angles, ratios, and distances. We also introduce a significantly faster and more robust symbolic engine, incorporating optimizations such as a reduced rule set and enhanced handling of double points. AG2 leverages a more powerful language model based on Gemini Team Gemini (2024) and trained on an order of magnitude larger and more diverse dataset. To further improve performance, we develop a novel search algorithm, Shared Knowledge Ensemble of Search Trees (SKEST), that explores a broader range of auxiliary construction strategies and employs a knowledge-sharing mechanism to expand and accelerate the search process. Finally, we make progress towards building a fully automated and reliable system that solves geometry problems in natural language. To do this we utilize Gemini to translate problems from natural language into the AlphaGeometry language and implement a new automated diagram generation algorithm.
@@
These enhancements culminate in a substantial improvement in performance. Specifically, AG2 achieves a new state-of-the-art solving rate of 84% on all IMO geometry problems from 2000 to 2024, compared to 54% achieved in AG1. This demonstrates a significant leap forward in AI’s ability to tackle challenging mathematical reasoning tasks, surpassing an average IMO gold medalist. In summary, the key improvements in AlphaGeometry2 include:
@@
  • Expanded domain language: covering locus-type theorems of moving objects, linear equations of angles/ratios/distances, and non-constructive problem statements.
@@
  • Stronger and faster symbolic engine: introducing an optimized rule set, better handling of double points, and a two-order of magnitude faster implementation in C++.
@@
  • Enhanced Language Model: leveraging the Gemini architecture, trained on an order of magnitude larger and more diverse dataset.
  • Advanced Novel Search Algorithm: introducing Shared Knowledge Ensemble of Search Trees (SKEST) that enables utilizing multiple search trees with knowledge sharing.
@@ 3 Automated problem formalization and diagram generation
Automating this process, called formalization, is an active area of research (see for example, Szegedy (2020); Wu et al. (2022); Jiang et al. (2023); Poiroux et al. (2024)). It is a significantly more complicated problem compared to a translation between human languages. While translation aims to preserve meaning, formalization frequently requires re-formulating the original problem into an alternative form, and sometimes disambiguating the nuances in the original problem statement. Automated formalization (auto-formalization), therefore, demands significant background knowledge and problem-solving skills on its own. Given that recently foundation models started demonstrating such capabilities, we use one such model, Gemini Team Gemini (2024), to automate the problem formalization for AlphaGeometry. We start by manually translating several dozens of geometry problems into the AG language. Then we use these examples to write a few-shot prompt asking Gemini to translate a given geometry problem from natural language into the AG language. We query Gemini five times with this prompt followed by another Gemini call asking to combine these results into one final answer. With this approach we are able to formalize 33 out of 44 formalizable IMO 2000-2024 geometry problems. For easier geometry problems, it is very consistent and makes almost no mistakes.
@@
Another manual part of our pipeline was diagram generation. In AG1, each point is defined by at most two basic predicates recalled in Table 1, the problem is therefore defined constructively and diagrams can be generated automatically. In AG2, we allow one or multiple points being defined simultaneously by an arbitrary number of predicates, allowing us to also cover non-constructive problems. Consider a non-constructive problem statement, “Let $ABC$ be a triangle with incenter $I$, such that $IA=2IB$”, here point $I$ is not only defined as an incenter, i.e. the intersection of two internal bisectors, but also defined by a third predicate $IA=2IB$ and there is no general strategy to construct such four points. Since AG2 covers non-constructive problems, diagram construction becomes a non-trivial part of the pipeline and generally requires human intervention. Similar to Krueger et al. (2021), below we propose a new algorithm to automatically generate diagrams given non-constructive problem specifications.
We start by initializing the points. We have three initialization methods which we alternate in different attempts:
  1. 1.
    Random – the points are chosen randomly using the normal distribution.
  2. 2.
    Construction in order.
  3. 3.
    Construction in a heuristically chosen order.
In order to construct a new point $X$, we look at each predicate $p(X,A_{1},\ldots,A_{k})$ involving $X$ point, and already constructed points $A_{1},\ldots,A_{k}$, and find a circle or line that X must be on. For example, if we know $|A_{1}A_{2}|=|A_{3}X|$, we have already constructed, we know $X$ should be on a circle with center $A_{3}$, and radius $|A_{1}A_{2}|$. If we cannot obtain such a circle, or a line, we skip that predicate. If there is not such set, $X$ is sampled randomly with a normal distribution. If there is exactly one such set, $X$ is sampled from it. If there are more such sets, $X$ is sampled to be any intersection of any two of them. We sample 10 examples for $X$, and choose one which minimizes the loss described below (with added noise to improve exploration).
After point initialization, we optimize their coordinates. Let $\bar{x}\in\mathbb{R}^{2n}$ be a vector representing all coordinates of all points. We encode every exact constraint $c$ in the diagram, including the goal, as $f_{c}(\bar{x})=0$ with a nonlinear function $f_{c}$, and each topological constraint as $g_{c}(\bar{x})<0$, or $h_{c}(\bar{x})\neq 0$. We numerically search for a suitable $\bar{x}$ in two steps. The overall loss for the Adam optimizer consists of
  1. 1.
    $\sum_{c\in C}f_{c}(\bar{x})^{2}$ where $C$ is the set of all the exact constraints,
  2. 2.
    $\sum_{c\in C}{\rm softplus}(g_{c}(\bar{x}))$ where $C$ is the set of all the topological constraints requiring $g_{c}(\bar{x})<0$,
  3. 3.
    $\sum_{c\in C}{\rm softplus}(\min(h_{c}(\bar{x}),-h_{c}(\bar{x}))$ where $C$ is the set of all the topological constraints requiring $h_{c}(\bar{x})=0$,
  4. 4.
    non-degeneracy loss of the form of a fraction of $L_{2}$ norm of all the point coordinates,
  5. 5.
    non-degeneracy loss of the form of $1/{(|AB|^{2}+\epsilon)}$ for all pairs of distinct points $A,B$.
We run the ADAM gradient descent optimization for a fixed number of steps simultaneously for 10 initial setups. Afterwards, we filter only the diagrams where the loss got below a given threshold, and all the topological constraints are satisfied. Finally, we switch from a gradient descent optimization to the Gauss-Newton-Levenberg method to look for a numerical solution of a combined under- and over-determined system of nonlinear equations.
@@
This three-stage optimization method is similar to the methodology introduced in Krueger et al. (2021). The final stage addresses the practical limitations encountered when tuning the gradient descent optimization in the original method, where achieving a consistently satisfactory error margin proved challenging.
@@
We benchmark this method on 44 IMO problems formalized in AG language (see Figure 8) and are able to find diagrams for 43. We run the three-stage convergence procedure in a loop which restarts and generates another random initial configuration after a failure. This way, 43 / 44 problems got their diagram sequentially generated within 1 hour.
@@ 9 Conclusions and Future work
Acknowledgement
Special thanks to Dawsen Hwang, Edward Lockhart, and Steven Creech for contributions to the development of AlphaGeometry2. We would also like to thank Yifeng Lu, Henryk Michalewski, Ed Chi, David Silver, Pushmeet Kohli, and Demis Hassabis for their thoughtful discussions, help and support.